Circular No. 7190

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999cm IN UGC 9766 T. Puckett and I. Siegert, Mountain Town, GA, report the discovery

of an apparent supernova (mag 17.1) on an unfiltered CCD frame (limiting mag 20.5) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on June 5.48 UT. SN 1999cm is located at α = 15^{h}11^{m}19^{s}.18, δ = +61^{o}07′11′′.4 (equinox 2000.0), which is 1′′.0 east and 15′′.1 south of the center of UGC 9766. The new object was also present on an unfiltered CCD frame taken on June 6.49, but it does not appear on a Palomar Sky Survey image taken on 1953 June 15 (limiting mag ∼ 20.9).

SUPERNOVA 1999cl IN NGC 4501 P. Garnavich, S. Jha, R. Kirshner, and P. Challis, Harvard-Smithson-

ian Center for Astrophysics, write: “Spectra obtained by M. Calkins with the 1.5-m Tillinghast telescope on June 4.2 UT show that SN 1999cl is a type-Ia supernova at or before maximum. A deep Si II absorption is seen with a minimum at 609.2 nm. Narrow emission lines from the host galaxy give a redshift of 2030 km/s, slightly smaller than the H I velocity of 2280 km/s (De Vaucouleurs et al. 1991, Third Reference Catalog of Bright Galaxies). The photospheric velocity derived from Si II is 14 500 km/s, which is quite high and implies that either the supernova is > 1 week before maximum or peculiar. An automatic type-Ia spectral aging algorithm (Riess et al. 1997, A.J. 114, 722) applied to our spectrum gives an age of −0.5 day. The absorption lines are particularly broad. The Si II absorption is observed out to a maximum blueshift of 22 200 km/s, and the S II lines that normally form a distinct ‘W’ at 533 nm are blended together. The continuum declines sharply toward the blue and, combined with narrow Na I absorption lines (host galaxy’s equivalent width = 0.33 nm; Galactic EW = 0.07 nm), this suggests that the supernova is heavily extinguished by dust. The spectral slope of SN 1999cl compared to SN 1998aq at an early phase gives an approximate color excess of E(B − V) = 1.0.”

COMET C/1999 H1 (LEE) Corrigendum. On IAUC 7183, line 7, for 1.08 ± 0.02 K km/s within

read 1.78 ± 0.06 K km/s (peak antenna temperature 1.08 ± 0.02 K), uncorrected for the efficiency of the antenna, within

1999 June 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green